### v1

@@ 4.2 Exp-2: FastER vs Batch-Query Baseline
Table 2 compares FastER and batch-query baselines across multiple datasets, focusing on recall and average time per matching result (Tavg). FastER consistently achieves the highest recall, demonstrating the effectiveness of rule filtering in capturing matching candidates. Regarding time efficiency, batch-query baselines require full dataset traversal, whereas our approach significantly enhances practical performance. Tavg comparisons show FastER achieves substantial speed improvements across most datasets, except for Amazon-Google.The reason for this exception lies in the large intra-cluster variance of the dataset, which makes it difficult for the rules to accurately capture all matching candidates. Attempting to capture all candidates would come at the cost of introducing a large number of irrelevant matches. Therefore, users need to carefully balance and weigh their decisions when mining and selecting rules. Specific optimization suggestions will be discussed in detail in Section 4.5.
@@ 4.3 Exp-3: FastER vs Progressive-Query Baselines
[…]
Fig. 6 shows error rate variations across datasets and $k$ values. FastER, as a precise method, consistently maintains a 0% error rate by verifying matches individually. In contrast, PBaseline exhibits high error rates in most datasets, except for SIGMOD20, which has minimal intra-cluster variance. Thus, PBaseline is unreliable for ER-on-demand.
Table 2: FastER vs Batch-Query baseline
Dataset Recall Tavg (seconds)
HG RobEM Ditto FastER HG RobEM Ditto FastER
FZ 0.954 1.000 1.000 1.000 2.781k 4.349k 0.917k 0.553k
DA 0.982 0.982 0.978 1.000 1.440k 3.198k 0.464k 0.052k
AG 0.654 0.846 0.786 0.841 1.162k 5.245k 0.818k 2.910k
WWC 0.902 0.585 0.927 0.923 1.069k 4.270k 1.105k 0.630k
GDS 0.943 0.800 0.900 1.000 1.817k 4.665k 0.678k 0.445k
Figure 6: FastER vs progressive-query baseline (error rate)
Table 3: Ablation Study of the FastER Framework
Method Arxiv CiteSeer WWC
Relative Comp. Recall Relative Comp. Recall Relative Comp. Recall
No-RF 453508.8 99.67% 14659.1 99.20% 117.9 99.00%
No-B 10.4 98.96% 29.4 98.67% 3.0 89.90%
No-PPS 1.6 97.78% 1.4 97.09% 2.3 89.20%
No-T 2.9 97.68% 3.0 96.42% 2.2 89.20%
FastER 1.0 97.68% 1.0 96.42% 1.0 89.20%
Method Fodors-Zagat DBLP-ACM Amazon-Google
Relative Comp. Recall Relative Comp. Recall Relative Comp. Recall
No-RF 1210.2 1 11662.8 1 301.1 99%
No-B 8.2 1 3.1 1 6.7 87%
No-PPS 1.2 1 1.6 1 1.4 82.18%
No-T 2.2 1 1.5 1 1.8 79%
FastER 1.0 1 1.0 1 1.0 79%
@@ 4.5 Exp-5: Rules Selection Strategies for Efficient Entity Resolution
When selecting a graph pattern, overly strict constraints should be avoided. For example: Pattern 1: $x\rightarrow y\rightarrow z\leftarrow y^{\prime}\leftarrow x^{\prime}$; Pattern 2: $x\rightarrow y\rightarrow z$, $x^{\prime}\rightarrow y^{\prime}\rightarrow z^{\prime}$. Requiring node $z^{\prime}$ to be strictly identical to $z$ may exclude many valid candidates. Instead, constraining the similarity between $z$ and $z^{\prime}$ rather than enforcing identical entity IDs (eid) prevents over-filtering and enhances match coverage. In summary, filtering rule and graph pattern selection directly influence the trade-off between execution time and recall. Stricter rules are preferable when execution time is a priority, whereas relatively looser rules are recommended when balancing execution time and recall. Furthermore, our latest research also explores an automated rule recommendation mechanism, and we will report further progress in future work (see GitHub).
@@ 5 Conclusion
We introduce FastER, an efficient on-demand entity resolution (ER) approach for property graphs based on Graph Differential Dependencies (GDDs), addressing the six key challenges outlined in Figure 1. FastER captures complex graph patterns for high-precision entity matching and employs Progressive Configuration Scheduling (PPS) to incrementally output results and provide real-time feedback. To enhance computational efficiency, FastER applies rule-based filtering, reducing candidate entity pairs by a factor of $10^{3}$ to $10^{5}$, and prioritizes user queries, focusing on user-specified subgraphs to avoid unnecessary computations. In dynamic environments, it recalculates only affected portions of the graph, preventing full-scale redundant processing, while pre-filtering matching functions to minimize computational overhead. Experimental results demonstrate that FastER significantly outperforms existing ER methods in both processing time and recall. Future research could explore advanced filtering techniques for heterogeneous data sources and automated rule selection mechanisms to better handle intra-cluster variations.
@@
Figure 8: Description of the GDD rules and its Cypher implementation.
Figure 9: Description of the GDD rules and its Cypher implementation.
Figure 10: Description of the GDD rules and its Cypher implementation.

### v2

@@ 4.2 Exp-1: FastER vs On-Demand Baselines
Table 2: FastER vs Batch-Query baseline
Dataset Recall Tavg (k seconds)
HG RobEM Ditto FastER HG RobEM Ditto FastER
[…]
@@ 4.3 Exp-2: FastER vs Batch-Query Baseline
Table 2 compares FastER and batch-query baselines across multiple datasets, focusing on recall and average time per matching result (Tavg). FastER consistently achieves the highest recall, demonstrating the effectiveness of rule filtering in capturing matching candidates. Regarding time efficiency, batch-query baselines require full dataset traversal, whereas our approach significantly enhances practical performance. Tavg comparisons show FastER achieves substantial speed improvements across most datasets, except for Amazon-Google. The reason for this exception lies in the large intra-cluster variance of the dataset, which makes it difficult for the rules to accurately capture all matching candidates. Attempting to capture all candidates would come at the cost of introducing a large number of irrelevant matches. Therefore, users need to carefully balance and weigh their decisions when mining and selecting rules. Specific optimization suggestions will be discussed in detail in subsection 4.6.
Figure 6: FastER vs Progressive-Query baseline (error rate)
Table 3: Ablation study results on graph and relational datasets.
Graph Datasets
Method Arxiv CiteSeer WWC
Rel. Comp. Recall (%) Rel. Comp. Recall (%) Rel. Comp. Recall (%)
No-RF 453 508.8453\,508.8 $99.67$ 14 659.114\,659.1 $99.20$ $117.9$ $99.00$
No-B $10.4$ $98.96$ $29.4$ $98.67$ $3.0$ $89.90$
No-PPS $1.6$ $97.78$ $1.4$ $97.09$ $2.3$ $89.20$
No-T $2.9$ $97.68$ $3.0$ $96.42$ $2.2$ $89.20$
FastER $1.0$ $97.68$ $1.0$ $96.42$ $1.0$ $89.20$
Relational Datasets
Method Fodors-Zagat DBLP-ACM Amazon-Google
Rel. Comp. Recall (%) Rel. Comp. Recall (%) Rel. Comp. Recall (%)
No-RF $1210.2$ $100.00$ 11 662.811\,662.8 $100.00$ $301.1$ $99.00$
No-B $8.2$ $100.00$ $3.1$ $100.00$ $6.7$ $87.00$
No-PPS $1.2$ $100.00$ $1.6$ $100.00$ $1.4$ $82.18$
No-T $2.2$ $100.00$ $1.5$ $100.00$ $1.8$ $79.00$
FastER $1.0$ $100.00$ $1.0$ $100.00$ $1.0$ $79.00$
@@ 4.4 Exp-3: FastER vs Progressive-Query Baselines
The most advanced progressive ER methods, such as PSN [11] and I-PES [9], rank entity pairs by matching likelihood, prioritizing the most probable matches to enhance early quality. However, new matches may alter the aggregated results, leading to inconsistencies. Unlike FastER, progressive baseline lacks a final verification mechanism to ensure strict on-demand output, contributing to its high error rate. We evaluate progressive baseline’s accuracy using the error rate over the first $k$ emitted entities, denoted as $\text{Err@}k$: $\text{Err@}k=\frac{\text{Errors}}{k}.$ Here, an error refers to an emitted result that does not meet user requirements, quantifying the proportion of incorrect matches in the first $k$ outputs.
Fig. 6 shows error rate variations across datasets and $k$ values. FastER, as a precise method, consistently maintains a 0% error rate by verifying matches individually. In contrast, progressive baseline exhibits high error rates in most datasets, except for SIGMOD20, which has minimal intra-cluster variance. Thus, progressive baseline is unreliable for ER-on-demand.
@@ 4.6 Exp-5: Rules Selection Strategies for Efficient Entity Resolution
When selecting a graph pattern, overly strict constraints should be avoided. For example: Pattern 1: $x\rightarrow y\rightarrow z\leftarrow y^{\prime}\leftarrow x^{\prime}$; Pattern 2: $x\rightarrow y\rightarrow z$, $x^{\prime}\rightarrow y^{\prime}\rightarrow z^{\prime}$. Requiring node $z^{\prime}$ to be strictly identical to $z$ may exclude many valid candidates. Instead, constraining the similarity between $z$ and $z^{\prime}$ rather than enforcing identical entity IDs (eid) prevents over-filtering and enhances match coverage. In summary, filtering rule and graph pattern selection directly influence the trade-off between execution time and recall. Stricter rules are preferable when execution time is a priority, whereas relatively looser rules are recommended when balancing execution time and recall. FastER is designed to leverage domain-specific rules for high-precision filtering. Meanwhile, recent work has begun to explore automated strategies for rule selection in graph-based entity resolution. We provide supporting resources on our project page [^2] https://github.com/Zaiwen/On_Demand_Entity_Resolution, which assist users in automatically constructing effective rules programmatically. This line of research is actively evolving and offers potential for future enhancements.
@@ 5 Conclusion
We introduce FastER, an efficient on-demand entity resolution (ER) approach for property graphs based on Graph Differential Dependencies (GDDs), addressing the six key challenges outlined in Figure 1. FastER captures complex graph patterns for high-precision entity matching and employs Progressive Profile Scheduling (PPS) to incrementally output results and provide real-time feedback. To enhance computational efficiency, FastER applies rule-based filtering, reducing candidate entity pairs by a factor of $10^{3}$ to $10^{5}$, and prioritizes user queries, focusing on user-specified subgraphs to avoid unnecessary computations. Experimental results demonstrate that FastER significantly outperforms existing ER methods in both processing time and recall. Future research could explore advanced filtering techniques for heterogeneous data sources and automated rule selection mechanisms to better handle intra-cluster variations.
Acknowledgment.
The authors affiliated with institutions $1,5$ and $6$ were supported in part by the National Key Research and Development Program of China under Grant No. 2023YFF1000100; the Hubei Key Research and Development Program of China under Grant Nos. 2024BBB055 and 2024BAA008; the Major Science and Technology Project of Yunnan Province under Grant No. 202502AE090003; and the Fundamental Research Funds for the Central Universities under Grant No. 2662025XXPY005.
Supplemental Material Statement.
Our supplementary material is available at https://doi.org/10.5281/zenodo.15389843 and includes the source code and datasets used for running FastER. All details are provided in the README.md file. Appendix.pdf contains the algorithm complexity proof and Cypher queries.
@@ Acknowledgment.
Figure 8: GDD rule examples with Cypher implementation (1).
Figure 9: GDD rule examples with Cypher implementation (2).
Figure 10: GDD rule examples with Cypher implementation (3).